A sample of argon gas takes up 7 L and has a temperature of 250 K. After being heated to temperature 280K, what is the new volume? (Use V₁/T₁ = V₂/T₂ ... Show the appropriate formula and math work for credit.)


1
Expert's answer
2021-06-11T22:47:53-0400

V1 = 7 L

T1 = 250 K

V2 = unknown

T2 = 280 K

P1 = P2 = const


Solution:

Since the pressure and amount of gas remain unchanged, Charles' Law can be used.

Charles' Law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the new volume, solve the equation for V2:

V2 = V1T2 / T1

V2 = (7 L × 280 K) / (250 K) = 7.84 L

V2 = 7.84 L


Answer: 7.84 L is the new volume.